OpenStudy (anonymous):

Find the value of the expression squar root of A^2 + 6^2 - 7^2 if a = 8 and b = 7

OpenStudy (anonymous):

where is b???

OpenStudy (anonymous):

is the 6 in the exp actually b?

OpenStudy (anonymous):

ya where is b

OpenStudy (anonymous):

soory i ment to put squar root of A^2 + 6^2 - B^2

OpenStudy (anonymous):

64+36-49

OpenStudy (anonymous):

just substitute the values in the expression then. what's the confusion?

OpenStudy (anonymous):

=51

OpenStudy (anonymous):

when i did the problem i got 51 then i sq and got 7.14 and its not an anser so i think i messed up some were

OpenStudy (anonymous):

\[a^{2}+6^{2}-b^{2}\] putting values \[8^{2}+6^{2}-7^{2}\] \[64+36-49\] = \[\sqrt {}\]

OpenStudy (anonymous):

=sqrt (51)

OpenStudy (anonymous):

what's the final answer given as?

OpenStudy (anonymous):

=8^2+6^2-7^2 =64+36-49 =651 take sq root of it =25.51 25.51 is the answer

OpenStudy (anonymous):

Check your solution once again @Qibtiya .

OpenStudy (jiteshmeghwal9):

put the given values in the equation as\[a^2+6^2-7^2\]\[8^2+6^2-7^2\]\[64+36-49=?\]